# Hausman’s Consistency Test and an Internal Bias Diagnostic: A Short Note Regarding Linear Mixed Models

### v1

@@ 5. Conclusion
We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects. Besides considering bias rather than consistency, this test is narrower than the Hausman test in the sense that it focuses on bias due to dependence between the random effects and their corresponding model matrix in situations where the fixed and random effects model matrices are not orthogonal (with respect to the inverse of the error covariance matrix). In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-$\boldsymbol{\eta}$ estimator of the fixed effects, provided the fixed-$\boldsymbol{\eta}$ estimator is consistent. Nevertheless, the bias diagnostic can rely solely on the fitted random-$\boldsymbol{\eta}$ model, generate summaries for each fixed effect or contrast, and use a permutation-based reference distribution. These features make it a potential complement to the Hausman test when assessing potential misspecification in linear mixed models.

### v2

@@ 5. Conclusion
We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects. Besides considering bias rather than consistency, this test is narrower than the Hausman test in the sense that it focuses on bias due to dependence between the random effects and their design matrix $\mathbf{Z}$} in situations where the fixed- and random-effects design matrices are not orthogonal (with respect to the inverse of the error covariance matrix). In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-effects estimator of the fixed effects, provided the fixed-effects estimator is consistent. Nevertheless, the bias diagnostic can be computed from the fitted random-effects model without changing the original RE specification and generates summaries for each fixed effect or contrast. These features make it a potential complement to the Hausman test when assessing potential misspecification in linear mixed models.
In applications where both RE and FE (or CRE) estimators are readily available, we view the Hausman or Mundlak–Wooldridge tests as primary specification checks: a rejection of the null typically leads the analyst to prefer an estimator that is robust to correlation between regressors and random effects. The bias diagnostic is then informative about which coefficients or contrasts are most affected and by how much. When the conclusions from the two approaches differ, we recommend using the diagnostic in a sensitivity-analysis role rather than as a replacement for the classical specification tests. A convenient workflow is therefore to (i) fit the RE model, (ii) apply a FE- or CRE-based Hausman/Mundlak–Wooldridge test, and (iii) when the test is rejected or borderline, use the mixed-model bias diagnostic to identify and report the coefficients or contrasts that are most affected.